OpenStudy (anonymous):

What is -5/6 ÷ 9 /10 ? A. -27/25 B. -4/3 C. -3/4 D. -25/27

OpenStudy (calculusxy):

Use \(\text{KCF}\). \(\text{K}\) keep the first fraction \(\text{C}\) change the operation sign to multiplication \(\text{F}\) flip (find the reciprocal) of the second fraction

OpenStudy (calculusxy):

\[\large -\frac{ 5 }{ 6 } \div \frac{ 9 }{ 10 } \rightarrow -\frac{ 5 }{ 6 } \times \frac{ 10 }{ 9 } = ?\]

OpenStudy (calculusxy):

Remember there is a negative.

OpenStudy (calculusxy):

\[\large -\frac{ 50 }{ 54 } \] \(\large \text{You can simplify this even more because there is a common factor.}\)

OpenStudy (calculusxy):

\(\text{You can simplify this even more because there is a common factor.}\)